OpenStudy (anonymous):

how do i simplify ( [3^(3+Δt)] - 3^3 )/Δt

OpenStudy (anonymous):

would changing 3^(3+Δt) to (3^3 )*(3^Δt) help?

OpenStudy (paxpolaris):

Δt=h \[=3^3{(3^h-1)\over h}\]

OpenStudy (anonymous):

what should i do next if there's a lim as h --> 0 in front of that?

OpenStudy (anonymous):

i need a way for all the h's to cancel out, i think

OpenStudy (paxpolaris):

you are supposed to use this method to find the derivative?

OpenStudy (anonymous):

yes

OpenStudy (paxpolaris):

i don't know how to show :\[\lim_{h \to 0} {3^h-1 \over h} = \ln(3)\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

the ln3 is what i forgot; i knew something equaled the first part of the equation!

OpenStudy (anonymous):

awesome, you're the best

OpenStudy (paxpolaris):

ok ... lol i am lost.